OpenStudy (anonymous):

The point P = (-3, 1, -1) is on a line, L, that is parallel to v = <2,4,-3>. Determine another point on L.

OpenStudy (amistre64):

just add 2 toghter

OpenStudy (anonymous):

L is parallel to V if its direction vector is equal?

OpenStudy (amistre64):

teh vector points to a position as well as a direction; so if you add it to the point; or even subtract it; or any scalar of it; you get a new point on the line

OpenStudy (amistre64):

By L and v being parallel; v is the vector for the line

OpenStudy (amistre64):

vector arent stationary, they move; just attach it to the point and count

OpenStudy (amistre64):

(-3, 1, -1) <2,4,-3> --------- -1,5,-4 is a point on the line

OpenStudy (amistre64):

in fact the line is comprised as: x=-3+2t y= 1+4t z=-1-3t

OpenStudy (anonymous):

Ah yea ok, since vectors are parallel if they have a scalar multiple.

OpenStudy (amistre64):

yep :)